# João Mattos, Debolina Halder Lina, Arlei Silva

### v1

@@ Abstract
Link prediction is a fundamental task in graph machine learning with applications, ranging from social recommendation to knowledge graph completion. Fairness in this setting is critical, as biased predictions can exacerbate societal inequalities. Prior work adopts a dyadic definition of fairness, enforcing fairness through demographic parity between intra-group and inter-group link predictions. However, we show that this dyadic framing can obscure underlying disparities across subgroups, allowing systemic biases to go undetected. Moreover, we argue that demographic parity does not meet desired properties for fairness assessment in ranking-based tasks such as link prediction. We formalize the limitations of existing fairness evaluations and propose a framework that enables a more expressive assessment. Additionally, we propose a lightweight post-processing method combined with decoupled link predictors that effectively mitigates bias and achieves state-of-the-art fairness–utility trade-offs.
@@ 1 Introduction
Link prediction is the task of discovering potential missing or future links in a network. There is extensive literature on link prediction models (liben2007link; zhang_link_2018; pan2021neural; zhu2021neural; chamberlain2022graph), which demonstrates the relevance of this task and the variety of domains where it can be applied, ranging from social networks to knowledge graphs. Among these domains, many types of networks are prone to biases in their structure and features (dai2024comprehensive; stoica2018algorithmic; karimi2018homophily), requiring the adoption of fair machine learning methods to mitigate bias in model predictions (li_dyadic_2021; li_fairlp_2022; current_fairegm_2022; tsioutsiouliklis2022link; tsioutsiouliklis2021fairness). Our work is focused on the fair link prediction problem and more specifically on the evaluation and design of link prediction models under fairness considerations.
@@
The graph representation learning literature has focused on fair node representations in the past (zhu_one_2024; zhu_devil_2024; ling_learning_2023; agarwal_towards_2021; laclau2021all), and building upon these works, the current notion of fairness adopted by fair link prediction methods is dyadic (li_dyadic_2021; current_fairegm_2022; li_fairlp_2022; luo2023cross). The main assumption behind the dyadic fairness framework is to categorize groups according to a protected attribute of interest and obtain equalized positive outcomes across these groups. In the social network example, two communities can be identified (e.g., men and women), and we can define fairness as links between men and women (inter) having the same probability of occurring as links within these groups (intra). To mitigate the disparity between these probabilities, fair link prediction algorithms adopt different strategies, evaluated through group fairness statistical metrics (dwork_fairness_2011; kusner_counterfactual_2017; masrour_bursting_2020).
@@
In addition, the dyadic fairness framework applied by previous fair link prediction methods (e.g. intra vs. inter) is unable to capture biases that occur within sensitive groups of node pairs. In the social network example, male-male connections might be systematically more likely than female-female connections and the dyadic fairness evaluation metrics considered (in particular, demographic parity) are insensitive to this phenomena, thereby masking underlying biases. The consequence of this limitation, is a ”glass ceiling” effect (stoica2018algorithmic), in which an under representation of a subgroup of pairs goes undetected.
@@
Limitations and Social Impact
Our evaluations are conducted on a limited set of graphs with small size, due to the unavailability of larger fairness datasets. Despite being flexible to non-binary sensitive attributes, our experiments are limited to binary sensitive attributes. We are not aware of any undesired societal impact caused by our work.
@@ 2 Preliminaries
Let a graph $G=(V,E,S)$, $v\in V$ is the set of nodes, $(u,v)\in E$ the set of existing edges in the graph, and $S\in\{0,1\}^{n}$ is the vector of sensitive attributes, where $s_{v}$ indicates the sensitive attribute of node $v$. In this work, we consider binary sensitive attributes for simplicity and availability of datasets, but our analysis is also valid for categorical and/or multiple sensitive attributes. The binary sensitive attribute $S$ produces three subgroups of node pairs, which for the remainder of the paper we denote as $E_{s\text{-}{s^{\prime}}}=\{(u,v)\in V\times V\mid s_{u}=1,s_{v}=0\}$, $E_{s\text{-}s}=\{(u,v)\in V\times V\mid s_{u}=1,s_{v}=1\}$, and $E_{s^{\prime}\text{-}s^{\prime}}=\{(u,v)\in V\times V\mid s_{u}=0,s_{v}=0\}$, for simplicity. We consider a link prediction classifier a score function $f(\cdot,\cdot)$ that maps a given input pair $(u,v)$ to a score $\hat{Y}\in[0,1]$. We denote the scores of all pairs candidate pairs as $R\in[0,1]^{|C|}$, where $C$ is the set of candidate pairs.
@@ 3 Limitations of Demographic Parity for Fair Link Prediction
Despite recent advances in fair graph learning, existing bias mitigation techniques for link prediction remain limited in three fundamental ways. First, many approaches rely on node representations from GNNs with constrained expressive power, restricting their ability to model the nuanced structural and demographic patterns necessary for fairness. Second, common fairness metrics such as demographic parity ($\Delta_{DP}$) operate under a dyadic assumption that aggregates across sensitive subgroups, masking imbalances that occur within sensitive groups. Finally, these metrics are often permutation-invariant and insensitive to ranking order, making them unsuitable for real-world applications where exposure and position matter. In this section, we systematically analyze these limitations and motivate the need for new fairness criteria tailored to link prediction.
@@ 3.1 Expressive Power Impact on Fairness
Most GNNs operate within the expressivity limits of the Weisfeiler-Lehman (1-WL) test, which restricts their ability to distinguish certain graph structures and node interactions (xu2018powerful). This poses a problem for fairness in link prediction, which often requires capturing subtle topological and demographic asymmetries across node pairs. For instance, a model must distinguish between link types (e.g., $E_{s^{\prime}\text{-}s^{\prime}}$ vs. $E_{s^{\prime}\text{-}s}$) and detect systematic underrepresentation. Yet, 1-WL GNNs tend to produce similar embeddings for nodes in symmetric neighborhoods, failing to differentiate pairwise combinations that matter for fairness. We include an example of this limitation in Appendix A.
@@
Figure 2: Illustration of how $\Delta_{DP}$ fails to capture exposure and subgroup proportion bias. Three models (a)–(c) output top-1010 link predictions over a graph with original subgroup edge proportions of 50% $E_{s\text{-}s}$ (blue), 30% $E_{s^{\prime}\text{-}s^{\prime}}$ (red), and 20% $E_{s^{\prime}\text{-}s}$ (black).
@@
Even if the final ranking distribution of pairs of each type in a ranking approximates the distribution from the original graph, $\Delta_{DP}$ is also limited by being permutation invariant. This characteristic also enables another instance of exposure bias (singh_fairness_2018), but this time in a dyadic fashion. For instance, a (biased) link prediction algorithm can output a ranking of pairs that promotes an unfair number of $E_{\text{intra}}$ ($E_{s\text{-}s}\bigcup E_{s^{\prime}\text{-}s^{\prime}}$) pairs to the top positions while still maintaining low values of demographic parity, characterizing a case of exposure disparity against $E_{\text{inter}}$. This is problematic in many applications, in which user attention is concentrated in the few top-ranked items.
@@
Figure 2 presents three hypothetical top-$10$ rankings produced by different link prediction models, each evaluated over the same input graph with subgroup edge distribution $\boldsymbol{\pi}=(0.5,0.2,0.3)$. Model (a) exhibits severe bias by overrepresenting $E_{s\text{-}s}$ pairs, distorting both the group proportions and their relative exposure in top ranks. Model (b) preserves the global edge distribution but places $E_{\text{intra}}$ edges disproportionately high, resulting in exposure bias despite matching overall proportions. Model (c) maintains both proportionality and fair exposure across groups. Despite these clear disparities, all three rankings could yield the same $\Delta_{DP}$ score, which is blind to ranking order and group-specific position effects—highlighting the need for rank-aware fairness evaluation.
@@
Previous definitions of exposure (singh_fairness_2018; zehlike_reducing_2020) consider the exposure of a group as the sum of the scores of items from the group weighted by the top-1 position bias. Such a definition assumes that items in top positions receive more attention (joachims_accurately_nodate) and, thus, should have larger weights associated with their scores. Following the same assumption, Property 2 ensures that significant deviations from the original graph distribution in the top positions are more heavily penalized than deviations occurring at the bottom positions.
@@ 4 Distribution-Preserving and Ranking-Aware Fair Link Prediction
In this section, we first demonstrate how adopting fair ranking information retrieval evaluation metrics for link prediction can address previous limitations associated with $\Delta_{DP}$ , and then we propose a simple post-processing algorithm for bias mitigation in link prediction methods.
@@ 4.1 Fair Ranking Metrics
The limitations of $\Delta_{DP}$ suggest the need for a fairness framework that accounts not only for global group proportions but also for exposure in ranked outputs. Inspired by previous works on fair ranking (zehlike_fair_2017), we pose the fair link prediction task as a group-aware ranking problem over candidate links.
@@ 4.2 MORAL - Multi-Output Ranking Aggregation for Link Fairness
At inference time, MORAL aggregates predictions from the group-specific models into a unified ranking. This is accomplished by maintaining a running estimate of the exposure distribution across the three edge types and greedily selecting, at each rank position, the highest-scoring remaining edge from the model whose inclusion most reduces the cumulative KL divergence from a predefined target distribution $\boldsymbol{\pi}$ (refer to Appendix F for pseudocode). This exposure-aware ranking procedure ensures that the final output approximates the desired group-wise exposure proportions.
Our greedy strategy solves the following fairness-prioritized objective: $\min_{R}\mathscr{L}_{A}(R)\quad\text{s.t.}\quad\mathscr{L}_{B}(R)\leq\min_{R^{\prime}}\mathscr{L}_{B}(R^{\prime})$. Similarly, one could optimize an objective that balances accuracy and fairness through a hyperparameter $\lambda$. A greedy algorithm can also achieve optimality in this weighted-sum objective by minimizing the combined loss at each step (celis2018ranking). We opt for the constrained formulation above to explicitly prioritize fairness across all datasets, especially in imbalanced settings where exposure risks are more severe.
@@ 5 Experiments
Table 1: Dataset statistics. Our analysis considers six real-world datasets. Results cover periphery and community graphs from different domains and varying sparsity levels.
@@
Baselines
To provide a robust comparison across different approaches for fair link prediction, we evaluate a diverse set of approaches and include the most competitive models. These include a node embedding method (FairWalk  (rahman_fairwalk_2019)); a pre-processing method (EDITS  (dong_edits_2022)); in-processing methods (GraphAIR  (ling_learning_2023) and UGE  (wang2022unbiased)); post-processing methods (DetConstSort (geyik_fairness-aware_2019) and MORAL); and a (task-specific) fair link prediction approach (FairAdj  (li_dyadic_2021)). For consistency, all methods are applied to a Graph Convolutional Network (GCN) encoder ($f_{\theta}$) with a dot product decoder, following the encoder-decoder framework outlined in Section 2. We expect similar results when considering other GNNs/SGNNs (wang_neural_2024; zhang_link_2018; chamberlain2022graph). We use the hyperparameter recommendations of the original papers in all experiments.
Datasets
We conduct experiments on six real-world datasets also considered in previous works (dong_edits_2022; li_fairlp_2022; current_fairegm_2022). These datasets represent diverse domains and fairness contexts. The Credit dataset is a credit scoring network where nodes represent individuals and edges represent credit relationships, with age as the sensitive attribute. The Facebook dataset is a social network where nodes are users and edges represent friendships, with gender as the sensitive attribute. The German dataset is a credit approval network where nodes are individuals and edges link similar individuals, with gender as the sensitive attribute. The NBA dataset is a network of NBA basketball players, where edges represent relationships between the athletes on Twitter, and nationality (’US’ and ’overseas’) is the sensitive attribute. The Pokec-n/z datasets consist of all the data from Pokec, a social network from Slovakia, in 2012, where nodes are users, edges are friendships, and gender is the sensitive attribute.
Table 1 presents statistics for each dataset, including the number of nodes, edges, and the distribution of sensitive attributes. We intentionally selected datasets from two different graph distributions (Community and Periphery).
Evaluation Metrics
We consider two metrics, adhering to the evaluation scheme of ranking tasks. We adopt NDKL as our fairness ranking metric and prec@k as our ranking-based performance metric. We also establish comparisons between NDKL and $\Delta_{DP}$ in Section 5.2.
Implementation Details
We implement all models using PyTorch Geometric (Fey/Lenssen/2019) and the PyGDebias (dong2023fairness) fairness library. We run experiments using NVIDIA A40 GPUs, fixing random seeds for reproducibility, and run each experiment 3 times. Random split edges are set as 70% for training, 10% for validation, and 20% for testing. We ensure that the sensitive attribute distribution is preserved across all splits. We adopt Adam optimizer with learning rate =0.0003=0.0003 for all experiments.

### v2

# Breaking the Dyadic Barrier: Rethinking Fairness in Link Prediction Beyond Demographic Parity

@@ Abstract
Link prediction is a fundamental task in graph machine learning with applications ranging from social recommendation to knowledge graph completion. Fairness in this setting is critical, as biased predictions can exacerbate societal inequalities. Prior work adopts a dyadic definition of fairness, enforcing fairness through demographic parity between intra-group and inter-group link predictions. However, we show that this dyadic framing can obscure underlying disparities across subgroups, allowing systemic biases to go undetected. Moreover, we argue that demographic parity does not meet the desired properties for fairness assessment in ranking-based tasks such as link prediction. We formalize the limitations of existing fairness evaluations and propose a framework that enables a more expressive assessment. Additionally, we propose a lightweight post-processing method combined with decoupled link predictors that effectively mitigates bias and achieves state-of-the-art fairness–utility trade-offs. [^1] This paper has been accepted for publication at the AAAI Conference on Artificial Intelligence (AAAI 2026).
@@ 1 Introduction
Link prediction is the task of discovering potential missing or future links in a network. There is extensive literature on link prediction models (liben2007link; zhang_link_2018; pan2021neural; zhu2021neural; chamberlain2022graph), which demonstrates the relevance of this task and the variety of domains where it can be applied, ranging from social networks to knowledge graphs. Among these domains, many types of networks are prone to biases in their structure and features (dai2024comprehensive; stoica2018algorithmic; karimi2018homophily), requiring the adoption of fair machine learning methods to mitigate bias in model predictions (li_dyadic_2021; li_fairlp_2022; current_fairegm_2022; tsioutsiouliklis2022link; tsioutsiouliklis2021fairness). Our work is focused on the fair link prediction problem and, more specifically, on the evaluation and design of link prediction models under fairness considerations.
@@
The graph representation learning literature has focused on fair node representations (zhu_one_2024; zhu_devil_2024; ling_learning_2023; agarwal_towards_2021; laclau2021all), and building upon these works, the current notion of fairness adopted by fair link prediction methods is dyadic (li_dyadic_2021; current_fairegm_2022; li_fairlp_2022; luo2023cross). The main assumption behind dyadic fairness is to categorize groups according to a protected attribute of interest and obtain equalized positive outcomes across these groups. In the social network example, two communities can be identified (e.g., men and women), and we can define fairness as links between men and women (inter) having the same probability of occurring as links within these groups (intra). To mitigate the disparity between these probabilities, fair link prediction algorithms adopt different strategies, evaluated through group fairness statistical metrics (dwork_fairness_2011; kusner_counterfactual_2017; masrour_bursting_2020).
@@
In addition, the dyadic fairness applied by previous fair link prediction methods (e.g., intra vs. inter) is unable to capture biases that occur within sensitive groups of node pairs, which is known as fairness gerrymandering (kearns2018preventing). In the social network example, male-male connections might be systematically more likely than female-female connections, and the dyadic fairness evaluation metrics considered (in particular, demographic parity) are insensitive to this phenomenon, thereby masking underlying biases. The consequence of this limitation is a ”glass ceiling” effect (stoica2018algorithmic), in which an under-representation of a subgroup of pairs goes undetected.
@@ 2 Preliminaries
Let a graph $G=(V,E,S)$, $v\in V$ is the set of nodes, $(u,v)\in E$ the set of existing edges in the graph, and $S\in\{0,1\}^{n}$ is the vector of sensitive attributes, where $s_{v}$ indicates the sensitive attribute of node $v$. In this work, we consider binary sensitive attributes for simplicity and the availability of datasets, but our analysis is also valid for categorical and/or multiple sensitive attributes. The binary sensitive attribute $S$ produces three subgroups of node pairs, which for the remainder of the paper we denote as $E_{s\text{-}{s^{\prime}}}=\{(u,v)\in V\times V\mid s_{u}=1,s_{v}=0\}$, $E_{s\text{-}s}=\{(u,v)\in V\times V\mid s_{u}=1,s_{v}=1\}$, and $E_{s^{\prime}\text{-}s^{\prime}}=\{(u,v)\in V\times V\mid s_{u}=0,s_{v}=0\}$, for simplicity. We consider a link prediction classifier a score function $f(\cdot,\cdot)$ that maps a given input pair $(u,v)$ to a score $\hat{Y}\in[0,1]$. We denote the scores of all candidate pairs as $R\in[0,1]^{|C|}$, where $C$ is the set of candidate pairs.
@@ 3 Limitations of Demographic Parity for Fair Link Prediction
Despite recent advances in fair graph learning, existing bias mitigation techniques for link prediction remain limited in three fundamental ways. First, many approaches rely on node representations from GNNs with constrained expressive power, restricting their ability to model the nuanced structural and demographic patterns necessary for fairness. Second, common fairness metrics such as demographic parity ($\Delta_{DP}$) operate under a dyadic assumption that aggregates across sensitive subgroups, masking imbalances that occur within groups. Finally, these metrics are often permutation-invariant and insensitive to ranking order, making them unsuitable for applications where exposure and position matter. We systematically analyze these limitations to motivate a new fairness criterion for link prediction.
@@ 3.1 Expressive Power Impact on Fairness
Most GNNs operate within the expressivity limits of the Weisfeiler-Lehman (1-WL) test, which restricts their ability to distinguish certain graph structures and node interactions (xu2018powerful). This poses a problem for fairness in link prediction, which often requires capturing subtle topological and demographic asymmetries across node pairs. For instance, a model must distinguish between link types (e.g., $E_{s^{\prime}\text{-}s^{\prime}}$ vs. $E_{s^{\prime}\text{-}s}$) and detect systematic under-representation. Yet, 1-WL GNNs tend to produce similar embeddings for nodes in symmetric neighborhoods, failing to differentiate pairwise combinations that matter for fairness. See Appendix A for an example of this limitation.
@@
Figure 2: $\Delta_{DP}$ fails to capture exposure and subgroup proportion bias. Three models (a)–(c) output top-1010 link predictions over a graph with original subgroup edge proportions of 50% $E_{s\text{-}s}$ (blue), 30% $E_{s^{\prime}\text{-}s^{\prime}}$ (red), and 20% $E_{s^{\prime}\text{-}s}$ (black).
@@
Even if the final ranking distribution of pairs of each type in a ranking approximates the distribution from the original graph, $\Delta_{DP}$ is also limited by being permutation invariant. This characteristic also enables another instance of exposure bias (singh_fairness_2018), but this time in a dyadic fashion. For instance, a (biased) link prediction algorithm can output a ranking of pairs that promotes an unfair number of $E_{\text{intra}}$ ($E_{s\text{-}s}\bigcup E_{s^{\prime}\text{-}s^{\prime}}$) pairs to the top positions while still maintaining low values of demographic parity, characterizing a case of exposure disparity against $E_{\text{inter}}$. This is problematic in many applications, in which user attention is concentrated on the few top-ranked items.
@@
Figure 2 presents three hypothetical top-$10$ rankings produced by different link prediction models, each evaluated over the same input graph with subgroup edge distribution $\boldsymbol{\pi}=(0.5,0.2,0.3)$. Model (a) exhibits severe bias by over-representing $E_{s\text{-}s}$ pairs, distorting both the group proportions and their relative exposure in top ranks. Model (b) preserves the global edge distribution but places $E_{\text{intra}}$ edges disproportionately high, resulting in exposure bias despite matching overall proportions. Model (c) maintains both proportionality and fair exposure across groups. Despite these clear disparities, all three rankings could yield the same $\Delta_{DP}$ score, which is blind to ranking order and group-specific position effects—highlighting the need for rank-aware fairness evaluation.
@@
Previous definitions of group exposure (singh_fairness_2018; zehlike_reducing_2020) consider the sum of the scores of items from the group weighted by the top-1 position bias. Such a definition assumes that items in top positions receive more attention (joachims_accurately_nodate) and, thus, should have larger weights associated with their scores. Following the same assumption, Property 2 ensures that significant deviations from the original graph distribution in the top positions are more heavily penalized than deviations occurring at the bottom positions.
@@ 4 Distribution-Preserving and Ranking-Aware Fair Link Prediction
We first demonstrate how fair ranking evaluation metrics can address previous limitations associated with $\Delta_{DP}$ , and then we propose a simple post-processing algorithm for bias mitigation in link prediction methods.
@@ 4.1 Fair Ranking Metrics
The limitations of $\Delta_{DP}$ suggest the need for accounting not only for group proportions but also for exposure in ranked outputs. Inspired by previous works on fair ranking (zehlike_fair_2017), we pose fair link prediction as a group-aware ranking problem over candidate links.
@@ 4.2 MORAL - Multi-Output Ranking Aggregation for Link Fairness
At inference time, MORAL aggregates predictions from the group-specific models into a unified ranking. This is accomplished by maintaining a running estimate of the exposure distribution across the three edge types and greedily selecting, at each rank position, the highest-scoring remaining edge from the model whose inclusion most reduces the cumulative KL divergence from a predefined target distribution $\boldsymbol{\pi}$ (see Algorithm 2 for pseudocode). This exposure-aware ranking procedure ensures that the final output approximates the desired group-wise exposure proportions.
Our greedy strategy solves the following fairness-prioritized objective: $\min_{R}\mathscr{L}_{A}(R)\quad\text{s.t.}\quad\mathscr{L}_{B}(R)\leq\min_{R^{\prime}}\mathscr{L}_{B}(R^{\prime})$. Similarly, one could optimize a different objective that balances accuracy and fairness through a hyperparameter $\lambda$. A greedy algorithm can optimize this weighted-sum objective by minimizing the combined loss at each step (celis2018ranking). We opt for the constrained formulation above to explicitly prioritize fairness across all datasets, especially in imbalanced settings where exposure risks are more severe.
@@
Input:
  • Target distribution $\boldsymbol{\pi}=(\pi_{0},\pi_{1},\pi_{2})$;
  • Total output size $n$.
Initialize exposure counts: $\boldsymbol{c}\leftarrow(0,0,0)$;
Initialize output ranking: $\mathbf{R}\leftarrow[\;]$;
for $t\leftarrow 1$ to $n$ do
 foreach group $j\in\{0,1,2\}$ such that $\mathcal{C}_{j}$ is not empty do
     Let $(u,v,\text{score})\leftarrow$ top element in $\mathcal{C}_{j}$;
    if this KL is lower than min_kl then
     end if
  end foreach
 Append $(\text{selected\_edge},\text{selected\_group})$ to $\mathbf{R}$;
 Remove top element from $\mathcal{C}_{\text{selected\_group}}$;
end for
return $\mathbf{R}$
@@ 5 Experiments
Table 1: Statistics from our six real-world datasets, covering periphery and community graphs from different domains.
@@
Baselines To enable a robust comparison for fair link prediction, we evaluate a diverse and competitive set of methods: a node embedding approach (FairWalk  (rahman_fairwalk_2019)); pre-processing methods (EDITS  (dong_edits_2022), FairLP  (li2022fairlp), FairEGM  (current2022fairegm)); in-processing methods (GraphAIR  (ling_learning_2023), UGE  (wang2022unbiased)); post-processing methods (DetConstSort  (geyik_fairness-aware_2019), MORAL); and a task-specific fair link prediction method (FairAdj  (li_dyadic_2021)). For consistency, all approaches use a GCN encoder ($f_{\theta}$) with a dot-product decoder, following the framework in Section 2. We expect similar trends with other GNNs/SGNNs (wang_neural_2024; zhang_link_2018; chamberlain2022graph). Hyperparameters follow each method’s original recommendations.
Datasets. We conduct experiments on six real-world datasets considered in previous works (dong_edits_2022; li_fairlp_2022; current_fairegm_2022). These datasets represent diverse domains and fairness contexts. The Credit dataset is a credit scoring network where nodes represent individuals and edges represent credit relationships, with age as sensitive attribute. The Facebook dataset is a social network where nodes are users and edges represent friendships, with gender as sensitive attribute. The German dataset is a credit approval network where nodes are individuals and edges link similar individuals, with gender as sensitive attribute. The NBA dataset is a network of NBA basketball players, where edges represent relationships between the athletes on Twitter, and nationality (’US’ and ’overseas’) is the sensitive attribute. The Pokec-n/z datasets consist of all the data from Pokec, a social network from Slovakia, in 2012, where nodes are users, edges are friendships, and gender is the sensitive attribute. Table 1 presents statistics for each dataset, including the number of nodes, edges, and the distribution of sensitive attributes. We selected datasets from two different graph distributions (community and periphery).
Evaluation Metrics. We consider two metrics, adhering to the evaluation scheme of ranking tasks. We adopt NDKL as our fairness ranking metric and prec@k as our ranking-based performance metric. We also establish comparisons between NDKL and $\Delta_{DP}$ in Section 5.2.
Implementation Details We implement all models using PyTorch Geometric (Fey/Lenssen/2019) and the PyGDebias (dong2023fairness). We run experiments using NVIDIA A40 GPUs, fixing random seeds, and run each experiment 3 times. Random edge splits are set as 70/10/20% for training/validation/testing. The sensitive attribute distribution is preserved across all splits. We adopt the Adam optimizer with learning rate $0.0003$ for all experiments.